If the ratio of the greater angle between diagonals of a rectangle to the 

smaller angle is 2: 1 and a > b then, the ratio a: b of sides of the rectangle is?


1
Expert's answer
2021-06-21T05:53:04-0400



Given "\varphi=2 (180^0-\varphi)"

"\varphi=360^0-2\varphi"

"3\varphi=360^0"

"\varphi=120^0"

"cot 30^0 = \frac{a}{b} \implies a:b = \sqrt{3}: 1"